OpenStudy (anonymous):

@ganeshie8 Please help me work out this problem step-by-step. I am confused. A gardener wants to divide a square piece of lawn in half diagonally. What is the length of the diagonal if the square is 8 ft? A. 4 B. 2[sqrt]8 C. 8[sqrt]2 D. 16[sqrt]8

OpenStudy (anonymous):

Is the length 16 if the diagonal is 8?

OpenStudy (anonymous):

I think it is 8[sqrt]2.

OpenStudy (anonymous):



OpenStudy (anonymous):

Note: it must be in simplest radical form.

OpenStudy (anonymous):

No the diagnol of a square or rectangle isnt quite as big as the 2 sides is splitting

OpenStudy (anonymous):

Area is 64. Is the diagonal 32?

OpenStudy (anonymous):

No thats not how it works here let me show you

OpenStudy (anonymous):

can someone answer my question?

OpenStudy (anonymous):

Sow you have 2 triangles the right angles sides both are 8 feet so what must the third "leg" be in length?

OpenStudy (anonymous):

I have given you the bases of your problem so try and solve it

OpenStudy (anonymous):

A^2+B^2=C^2

OpenStudy (anonymous):

8^2+8^2=64

OpenStudy (anonymous):

64^2=4096

OpenStudy (anonymous):

No your not finding the area you are finding the perimeter of a triangle

OpenStudy (anonymous):

P=32

OpenStudy (anonymous):

32-16=?

OpenStudy (anonymous):

16.

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

That is not one of my answer choices, no.

OpenStudy (anonymous):

I dont think the var is 16 in length of the right angle triangle now i gave you the parts and i gtg. So just use what i gave you i gave you all i could to help without cheating

OpenStudy (anonymous):

Fan me if you want me later

OpenStudy (anonymous):

Okay, I'm just going to make an educated guess. I don't understand.

OpenStudy (anonymous):

I will ask my teacher.